OpenStudy (mallorysipp234):

this makes no sense. please help. Using your exponential function, P(x), explain the meaning behind each of the numbers and variables in terms of the invasive fish species. my exponential function is 10 x 2^x

OpenStudy (anonymous):

a lot of stuff doesn't make sense.... lol :P

OpenStudy (mallorysipp234):

i understand how to describe the terms & variables but i dont get what it means by the "invasive fish species" part. like...what.

OpenStudy (anonymous):

is there stuff before this on the invasive fish species?

OpenStudy (mallorysipp234):

no.

OpenStudy (mallorysipp234):

that's why im confused 'af

OpenStudy (anonymous):

hmmm.... ok then

OpenStudy (mallorysipp234):

should i just write down what the terms and variables mean? i just did that for the previous question though.

OpenStudy (anonymous):

i found this online... A fisherman decided to stock his favorite lake with a few fish that are not native to the region. The fish he introduces have no natural predators in the lake and they experience a large exponential growth. The introduction of an invasive species can be very detrimental to an ecosystem and there have already been noted effects. The National Fish and Wildlife Foundation has asked for your help. 1) Create the data to fill in the population table. Be creative when selecting the P(x) values, but remember that this is an exponential growth model. Notice that you even need to come up with how many fish the fisherman introduces to the lake. Months (x) 0 1 2 3 P(x) 2) Describe how to create the exponential function represented by your table. Use complete sentences and arrive at a final exponential function, P(x). 3)Using your exponential function, P(x), explain the meaning behind each of the numbers and variables in terms of the invasive fish species. 4) The current exponential function P(x) measures the fish population’s growth monthly. Using complete sentences explain how to find the rate of growth every week. 5)The National Fish and Wildlife Foundation has asked you to calculate when the population of the fish reaches 500. Show all of your work. 6)The computer system you need to input the function into only works in logarithms of base 10. Using complete sentences, explain how to convert your exponential function P(x) in a logarithmic one and then into a base 10 logarithm.

OpenStudy (mallorysipp234):

yup. thats what im working on.

OpenStudy (anonymous):

OOOOOOOOOOOHHHHHHHHHH... OK... give me a sec then...

OpenStudy (mallorysipp234):

alright...

OpenStudy (anonymous):

10 x 2^x 10 is the coefficient on the variable... this means that there are 10x- 10 fish originally... i think. and I think the 2 is an exponent.... am i correct?

OpenStudy (anonymous):

or is the x an exponent?

OpenStudy (mallorysipp234):

p(x) = 10 x 2 ^ x

OpenStudy (mallorysipp234):

x is the small number on top. the exponent to the 2

OpenStudy (anonymous):

ok

OpenStudy (mallorysipp234):

\[p(x) 10 x 2^{x}\]

OpenStudy (anonymous):

so the x is the rate that the number of fish changes. this is unknown, seeing as it is a variable

OpenStudy (mallorysipp234):

i need help with number 5. how to tell when the fish population reaches 500

OpenStudy (anonymous):

so.... to sum it up... 10 - the original number of fish x - the rate the amount of fish changes

OpenStudy (anonymous):

ok... i can help

Parth (parthkohli):

x is the times sign, not a variable.

OpenStudy (anonymous):

it is?

OpenStudy (mallorysipp234):

yes

Parth (parthkohli):

yes, it is.

OpenStudy (mallorysipp234):

i need help with five and six.

OpenStudy (mallorysipp234):

please

OpenStudy (anonymous):

ok

Parth (parthkohli):

\[10 \times2^x\]you just use this equation to describe some kind of change.

OpenStudy (anonymous):

no... you would plug it in for y; p(x) is equal to y

OpenStudy (mallorysipp234):

this is for #5 right?

Parth (parthkohli):

probably the number of fish.

OpenStudy (anonymous):

y = 10 * 2 ^ x yeah

Parth (parthkohli):

\[10\times 2^ x = 500\]

OpenStudy (anonymous):

for me... at least....

OpenStudy (mallorysipp234):

okay but i need to figure out WHEN the fish population reaches 500.

OpenStudy (anonymous):

so then you plug in 500 for y and solve for x. 500 = 10 * 2 ^ x divide by 10... 50 = 2^x then you solve.... i got 5.64... this means 5.64 months

OpenStudy (anonymous):

does that help?

OpenStudy (mallorysipp234):

THANK YOU SO MUCH YES omg can you help me with number 6?

OpenStudy (anonymous):

sure :D

OpenStudy (anonymous):

Exponential growth functions are y=bx and the form of logs are y=log5 (x). So, in my function, 5 is b and x is x. So the log would be log_5 (x). To convert it to where it will be a log with a base of 10, I am going to substitute with 5 because that is how many fish are in the population at the end of 1 month. Okay so now we have log_5 (5). Now we can properly convert this. Ln(5)/ln(5)= 1. So we take that and plug it into log_10 (x) and get log_10 (1). And that is how we convert our exponential function into logs...

OpenStudy (anonymous):

make sense? or should i simplify it?

OpenStudy (mallorysipp234):

at the end of month 1 there are 20 fish in the pond not 5.

OpenStudy (mallorysipp234):

months population p(x) 0 10 1 20 2 40 3 80

OpenStudy (anonymous):

ohh.. oops...yeah ur right.... soooooo.... let me fix that.... Exponential growth functions are y=bx and the form of logs are y=log5 (x). So, in my function, 5 is b and x is x. So the log would be log_5 (x). To convert it to where it will be a log with a base of 10, I am going to substitute with 20 because that is how many fish are in the population at the end of 1 month. Okay so now we have log_5(20). Now we can properly convert this. Ln(20)/ln(20)= 1. So we take that and plug it into log_10 (x) and get log_10 (1). And that is how we convert our exponential function into logs.

OpenStudy (mallorysipp234):

alright thank you so much (: that all makes a lot more sense now.

OpenStudy (anonymous):

yeah! you are welcome :D